A body is dropped from the top of the tower falls 100m during last 2 sec of its falls. The time taken  to reach the ground from the top of the tower

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Let total time to reach the ground bet.

Hence distance travelled in time't'iss_t=\dfrac{1}{2}gt^2

Distance travelled in time ‘t-2’ iss_{t-2}=\dfrac{1}{2}g(t-2)^2

Therefore100=\dfrac{1}{2}g(t^2-(t-2)^2)
\implies t=6s